OpenStudy (solomonzelman):

Question about desmos calculator. Is it possible to graph a piece wise function on it?

OpenStudy (solomonzelman):

Like if I need f(x) = e^x if x<0 (i.e. (-∞,0) ) = x^2 if x≥0 (i.e. [0,∞) )

OpenStudy (anonymous):

For the first entry, type (exactly) \[y=e^x\{x<0\}\] and for the second entry, \[y=x^2\{x\ge0\}\] There's a button for the \(\ge\) symbol.

OpenStudy (anonymous):

You can also include constants like \(a,b,c\), etc. You'll get an option to include a slider that alters the values of the constants. Pretty neat.

OpenStudy (solomonzelman):

yeah, I can put up on the keyboard ≥ and ≤... thank you for the help !!